OpenStudy (anonymous):

At the city museum, child admission is $6.40 and adult admission is $9.90 . On Friday, 184 tickets were sold for a total sales of $1524.10 . How many adult tickets were sold that day?

OpenStudy (anonymous):

let x be the number of children let y be the number of adults x+y= 184 (eqn 1) 6.40x + 9.90y = 1524.10 (eqn 2) you are looking for the value of y or the number of adult tickets sold.. you can use substitution method in solving for y.. do you know how to do it? evaluate the first equation in terms of x first.. tell me what you got and i tell you what's the next step

OpenStudy (anonymous):

im so confused hahahaa

OpenStudy (anonymous):

which part are you confused with? i'll try to explain (:

OpenStudy (anonymous):

x=184-y

OpenStudy (anonymous):

yes that's right! since x=184-y replace the variable x in the second eqn using the new evaluated eqn.. like this: \(\sf 6.40 \color{red}{x} + 9.90y = 1524.10\\\) since \(\sf \color{red}{x=184-y}\), you'll have: \(\sf 6.40 (\color{red}{184-y}) + 9.90y = 1524.10\\\) can you simplify this and solve for y?

OpenStudy (anonymous):

oh ok ! is it 99 adult tickets ?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

np welcome to openstudy !